OpenStudy (anonymous):

A tank has 2 taps. If tap A is open, the tank drains empty in 10 min. If tap B is open the tank drains empty in 6 min. How long would it take to drain the tank if both taps are open?

OpenStudy (anonymous):

The way to think about these is to realize that one tap is faster, or more efficient, at draining. Say, for example, tap A drains 60 gallons per minute. The tank drains in 10 minutes, so there are 10 x 60 = 600 gallons. Using tap B, the 600 gallons drains in 6 minutes, so tank B must drain at a rate of 100 gallons per minute. It turns out that you don't have to know how many gallons are in the tank to get started... that's just an example.

OpenStudy (anonymous):

Another possibly easier way to set it up is to say: t = time to drain using both taps t x (drain rate of tap A) + t x (drain rate of tap B) = whole tank drained so, using my earlier example to make the math easy, let's assume there are 600 gallons in the tank: t (60 gallons per minute) + t (100 gallons per minute) = 600 gallons t (160) = 600 t = 600/160 = 3.75 minutes. As I said, it doesn't matter what the tank size is.

OpenStudy (anonymous):

t (tank/10 minutes) + t (tank/6 minutes) = tank all three terms in the above equation have "tank" as a variable, so you can cancel it out, leaving: t/10 + t/6 = 1 Solve for t... t will be = 3.75